# SLMRec: Empowering Small Language Models for Sequential Recommendation

### v1

@@ 5 Experiments
In this section, we present extensive experiments to demonstrate the effectiveness of $\mathrm{SLMRec}$, aiming to answer the following three research questions (RQs).
@@
$\bullet$ RQ4: Is it feasible to train our SLMRec model without utilizing a pre-trained teacher model?
@@ 5.1 Experiment Setup
Table 1: Statistics on the Amazon datasets.
@@
  • $\left|\mathcal{U}\right|$, $\left|\mathcal{V}\right|$, $\left|\mathcal{E}\right|$ denote the number of user, item and ratings, respectively.
For our experimental evaluation, we utilize data from the clothing, movies, music, and sports categories within the extensive, industry-scale Amazon18 dataset  [^3] https://nijianmo.github.io/amazon/index.html. Statistics of the datasets are shown in Table 1. In all datasets, we interpret any rating above 3 as implicit feedback, indicating user interaction with the item, and employ timestamps to establish the chronological order of actions. We eliminate users and items that have fewer than 5 associated actions to ensure sufficient data density. The historical sequence of interactions for each user is divided into three segments: (1) the most recent interaction is reserved for testing, (2) the second most recent for validation, and (3) all preceding interactions are used for training. Based on the ranking results, we utilize the typical $top$-$N$ metrics hit rate (HR@{1, 5, 10}), normalized discounted cumulative gain (NDCG@{5,10}) [26] and Mean Reciprocal Rank (MRR) [45] to evaluate the model performance. For all the metrics, higher values indicate better performance. Models that achieve the highest MRR performance on the validation set, including ours and other baseline models, will be preserved for subsequent performance evaluation on the test set. In order to ensure an unbiased evaluation, we adopt the methodology employed in previous works [33; 58], wherein we randomly select 999 negative items (i.e., items that the user has not interacted with) and combine them with 1 positive item (i.e., a ground-truth interaction) to form our recommendation candidates for the ranking test. Detailed hyperparameters of our model in each dataset can be obtained in Appendix A.1.
@@ 5.2 Performance Comparisons
Compared Methods. We compare our method with three classes of baselines: (1) Single-domain sequential recommendation methods, i.e., BERT4Rec [46], GRU4Rec [18] and SASRec [29]. (2) G-LLMRec method: Open-P5LLaMa [^4] For Open-P5, we adopt the version of LLaMa as the foundation model in their code repository implementation to ensure the best results are achieved. [54]. (3) E-LLMRec method: E4SRec* [35]. A detailed introduction to these baselines can be found in Appendix A.2. It should be noted that we did not select various G-LLMRec methods or E-LLMRec methods as baselines. This is because the differences between each LLM-based method are minimal, and our model is a universal approach that is not confined to a specific model type. Our primary focus is to improve the efficiency of language model utilization. Hence, we opted to select one generative-based method (Open-P5) and one embedding-based method (E4SRec) as our baselines.
Table 2: Experimental results (%) on the Cloth and Movie dataset. The missing MRR value of Open-P5 is unavailable due to the time complexity constrictions. The number on the left of the arrow is the layers $N$ of the student model. The left number on the right of the arrow is the layers $M$ of the teacher model.
Methods Cloth Dataset Movie Dataset Average
Improv.
HR NDCG MRR HR NDCG MRR
@1 @5 @10 @5 @10 @1 @5 @10 @5 @10
GRU4Rec [18] 13.79 15.46 16.83 14.64 15.08 15.15 10.56 19.47 25.21 15.11 16.96 15.46 -20.26
BERT4Rec [46] 13.60 14.66 15.55 14.14 14.43 14.59 9.68 14.91 17.98 12.40 13.38 12.74 -29.84
SASRec [29] 13.08 16.94 20.26 15.01 16.08 15.76 5.57 16.80 26.85 11.17 14.42 12.08 -25.19
Open-P5LLaMa [54] 14.13 17.68 19.74 17.02 16.40 - 12.66 21.98 27.24 17.13 19.81 - -9.96
E4SRec* [35] 16.71 19.45 21.86 18.09 18.86 18.77 14.74 23.79 29.09 19.45 21.16 19.74 0.00
E4SRec8[35] 15.30 18.54 21.29 16.91 17.79 17.60 13.32 22.49 28.57 17.99 19.94 18.46 -5.90
E4SRec4[35] 14.58 18.05 20.92 16.32 17.25 17.01 11.80 21.54 28.02 16.73 18.82 17.20 -10.24
$\mathrm{SLMRec}_{8\leftarrow 32}$ 16.56 19.05 21.33 17.79 18.53 18.48 15.18 23.93 29.30 19.69 21.41 20.06 -0.17
$\mathrm{SLMRec}_{4\leftarrow 32}$ 14.86 18.03 20.70 16.45 17.30 17.12 13.70 22.73 28.44 18.37 20.21 18.74 -6.56
$\mathrm{SLMRec}_{4\leftarrow 8}$ 16.10 18.85 21.33 17.48 18.28 18.17 14.83 23.08 28.02 19.08 20.67 19.45 -2.55
+$\mathcal{D}_{norm}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 16.28 19.12 21.75 17.69 18.53 18.40 14.86 23.89 30.22 19.36 21.39 19.84 -0.37
+$\mathcal{L}_{ms}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 16.85 19.05 20.93 17.96 18.57 18.59 15.05 23.48 28.60 19.40 21.05 19.76 -0.85
+$\mathcal{D}_{norm}$+$\mathcal{L}_{ms}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 16.69 19.47 21.90 18.07 18.85 18.74 15.29 24.25 30.19 19.90 21.82 20.36 +1.49
@@
Methods Music Dataset Sport Dataset Average
Improv.
HR NDCG MRR HR NDCG MRR
@1 @5 @10 @5 @10 @1 @5 @10 @5 @10
GRU4Rec [18] 1.89 3.22 4.36 2.57 2.93 3.08 5.26 7.75 9.64 6.52 7.13 7.08 -41.93
BERT4Rec [46] 2.10 3.16 4.04 2.64 2.92 3.11 4.81 6.70 7.98 5.79 6.20 6.26 -46.82
SASRec [29] 1.82 5.72 8.64 3.79 4.72 4.51 4.70 8.43 11.21 6.59 7.48 7.24 -30.60
Open-P5LLaMa [54] 4.35 8.12 10.18 6.74 7.83 - 5.49 8.5 11.28 6.92 7.65 - -12.62
E4SRec* [35] 5.62 9.29 11.64 7.50 8.26 7.98 6.40 9.67 12.35 8.05 8.91 8.70 0.00
E4SRec8[35] 5.46 8.86 11.14 7.21 7.95 7.74 5.48 8.63 11.34 7.06 7.93 7.76 -7.48
E4SRec4[35] 5.33 8.75 10.94 7.08 7.78 7.59 5.41 8.65 11.35 7.04 7.91 7.72 -8.54
$\mathrm{SLMRec}_{8\leftarrow 32}$ 5.65 8.89 11.34 7.30 8.09 7.90 5.43 8.77 11.45 7.11 7.97 7.78 -6.39
$\mathrm{SLMRec}_{4\leftarrow 32}$ 5.26 8.55 10.88 6.96 7.71 7.50 5.76 8.91 11.45 7.35 8.17 8.02 -7.56
$\mathrm{SLMRec}_{4\leftarrow 8}$ 5.62 8.78 11.09 7.23 7.97 7.81 6.25 9.25 11.68 7.76 8.54 8.41 -3.55
+$\mathcal{D}_{norm}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 5.95 9.26 11.68 7.65 8.42 8.23 6.61 9.82 12.34 8.24 9.05 8.87 +1.97
+$\mathcal{L}_{ms}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 5.69 8.94 11.18 7.36 8.08 7.91 6.51 9.39 11.80 7.96 8.74 8.62 -1.75
+$\mathcal{D}_{norm}$+$\mathcal{L}_{ms}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 5.72 9.15 11.45 7.48 8.21 8.03 6.62 9.83 12.34 8.25 9.06 8.89 +0.81
@@
Quantitative Results (RQ1). Tables 23 showcase the quantitative comparison of four large-scale sequential recommendation datasets. From our analysis, we have several insightful observations: (1) LLM-based recommendation methods exhibit substantial improvements over traditional sequential recommendation (TSR) methods, primarily due to their enhanced modeling capacity which adeptly extracts informative sequential interest patterns. (2) Our model, $\mathrm{SLMRec}_{4\leftarrow 8}$, outperforms the teacher model E4SRec8* by leveraging knowledge distillation within the hidden layers. By refraining from applying this constraint prior to the prediction phase, we enable the final representation to organically gravitate towards the label—yielding an approximate 8% enhancement in performance in comparison to the teacher model. (3) Introducing vanilla knowledge distillation techniques into LLMRec, without altering the model structure, allows $\mathrm{SLMRec}_{4\leftarrow 8}$ to achieve a marginally superior performance compared to E4SRec32*. This suggests that small language models equipped with efficacious training strategies can rival, or even exceed, larger language models in the sequential recommendation task.
@@
Model Efficiency (RQ2). We report the time efficiency and parameters of comparative baselines and our model in Table 4. All time and parameter metrics represent the average across the four datasets reported. Inference time evaluates the prediction ranking among 1,000 candidate items for each user. Detailed training and inference times for each dataset are provided in Appendix A.3. The Open-P5, an LLMRec model based on generative methods, offers a reasonable training duration. Yet, during the inference phase, it becomes considerably time-consuming (4942 hours) as it necessitates generating a substantial pool of candidate items (for instance, 1000). Owing to the intrinsic workings of generative LLMs, employing generation-based LLMRec models for the comprehensive ranking of extensive item sets is not advised. Our model outperforms E4SRec with enhanced efficiency, maintaining only 13% and 14% in E4SRec’s parameters for training and inference, respectively. Moreover, our SLMRec demonstrates a remarkable gain in speed, being 6.6 times faster during training and 8.0 times quicker in inference than E4SRec.
@@
Ablation Study (RQ3). In all experimental results, SLMRec, when enhanced with various knowledge regularizers (namely $\mathcal{D}_{cos},\mathcal{D}_{norm}$ and $\mathcal{L}_{ms}$), demonstrates improved performance. The regularizers $\mathcal{D}_{cos}$ and $\mathcal{D}_{norm}$ aid SLMRec in aligning its intermediate representations with those of the teacher model, thereby endowing it with more potent representational extraction capabilities. Meanwhile, $\mathcal{L}_{ms}$ steers the model to assimilate domain knowledge pertinent to recommendation systems within its preliminary layers.
@@ 5.3 Model Study
Figure 4: Model study of online KD and block number $b$ in the music dataset.
@@ 6 Related Work
In this section, we introduce the most related background and scientific investigations to this work, which are roughly divided into five categories, i.e., 1) Sequential Recommendation, 2) Knowledge Distillation (KD) 3) Depth-wise Knowledge of LLMs, 4) Model Pruning, and 5) Parameter-Efficient Fine-Tuning (PEFT).
@@
Sequential Recommendation. Traditional Sequential Recommendation (TSR) methods [52; 18; 46; 29] primarily focus on developing various temporal encoders to capture short- and long-term user interests. Following the triumph of large language models (LLMs), researchers have begun leveraging open-source LLMs [47] to construct their recommendation systems [27; 5; 51]. G-LLMRec methods [12; 54; 57; 38] generate the next item based on historical sequences, while E-LLMRec approaches [35; 59; 50] use LLMs as feature extractors to learn user representations for prediction. More recently,  [55] introduces a generative sequential framework scalable up to GPT-3 dimensions, trained from scratch with 100 billion examples and utilizing 256 H100 GPUs. LLM-based recommendation systems frequently outperform TSR models by a margin of 20% [35; 38; 50], also increasing the parameters by nearly 100 times compared to TSR models. Therefore, the deployment of LLMRec models in real-world platforms is heavily constrained by computational resources.
@@
Knowledge Distillation (KD). Training a smaller “student" model on the distribution predicted by a large “teacher" model is known as a powerful knowledge distillation technique [19]. The fundamental insight behind this is to transform the knowledge and capabilities of the teacher into more compact, compressed, and possibly skill-specific representations [28; 15]. For those cases when the student only has access to the output tokens generated by the teacher, another way of KD is data distillation [9; 36; 11; 23]. This technique first generates high-quality synthetic data by prompting the larger teacher model. The synthetic data are then used to enhance the student’s capabilities by fine-tuning. Our work lies in the former series, which performs a simple yet effective layer-skipping approach to transfer those useful representations of the teacher to the student.
Depth-wise Knowledge of LLMs. The recent community interest stems from how linguistic properties and knowledge are encoded in language models.  [41; 7] emphasize that knowledge localizes within the middle or final layers. On the other hand,  [16] attempts to perform knowledge editing and concludes that information may be stored non-locally across layers. What’s more,  [40; 14] share a similar view that current pretraining methods are not properly leveraging the parameters in the deeper layers of the network or that the shallow layers play a critical role in storing knowledge. By contrast, we are the first to investigate which part of knowledge on the LLMs plays a key role, especially in the sequential recommendation scene.
Model Pruning. Model Pruning is a fundamental approach for reducing the size of a well-trained large model by removing unimportant parameters [17]. Recent work has focused on applying pruning methods to the Transformer architecture [48]. These works have studied different components of the model architecture for pruning, including dropping attention heads [49; 42], dropping layers  [10; 56; 31; 44], dropping hidden states [21], replacing sparse weight matrices with smaller dense ones [4], and combinations of these solutions. By contrast, our work performs layer removal through simple knowledge distillation, rather than more complex pruning techniques.
Parameter-Efficient Fine-Tuning (PEFT). PEFT emerges as a novel technique for tailoring Large Language Models (LLMs) to specific tasks while ensuring minimal computational and memory costs [22; 34; 25; 39]. In this work, we combine our method with the Low-Rank Adapters (LoRA) [25] to reduce the memory and computation of the knowledge distillation process. Specifically, we freeze the pre-trained model and only tune a small set of additional trainable parameters.
7 Conclusions and Outlooks

### v3

@@ 5 Experiments
In this section, we present extensive experiments to demonstrate the effectiveness of SLMRec, aiming to answer the following four research questions (RQs).
@@
$\bullet$ RQ4: Is it feasible to train our model, SLMRec, simultaneously with an untrained teacher model?
@@ 5.1 Experiment Setup
Table 1: Statistics of the Amazon datasets. $\left|\mathcal{U}\right|$, $\left|\mathcal{V}\right|$, and $\left|\mathcal{E}\right|$ denote the number of users, items, and ratings, respectively.
@@
For our experimental evaluation, we utilize data from the clothing, movies, music, and sports categories within the extensive, industry-scale Amazon18 dataset [^4] https://nijianmo.github.io/amazon/index.html. Statistics of the datasets are shown in Table 1. In all datasets, we interpret any rating above 3 as positive feedback, indicating user interaction with the item, and employ timestamps to establish the chronological order of actions. We eliminate users and items that have fewer than 5 associated actions to ensure sufficient data density. The historical sequence of interactions for each user is divided into three segments: (1) the most recent interaction is reserved for testing, (2) the second most recent for validation, and (3) all preceding interactions are used for training. Based on the ranking results, we utilize the typical $top$-$N$ metrics hit rate (HR@{1, 5, 10}), normalized discounted cumulative gain (NDCG@{5,10}) (Järvelin & Kekäläinen, 2002) and Mean Reciprocal Rank (MRR) (Sarwar et al., 2001) to evaluate the model performance. For all the metrics, higher values indicate better performance. Models that achieve the highest MRR performance on the validation set, including ours and other baseline models, will be preserved for subsequent performance evaluation on the test set. In order to ensure an unbiased evaluation, we adopt the methodology employed in previous works (Krichene & Rendle, 2020; Zhao et al., 2020), wherein we randomly select 999 negative items (i.e., items that the user has not interacted with) and combine them with 1 positive item (i.e., a ground-truth interaction) to form our recommendation candidates for the ranking test. Detailed hyperparameters of our model in each dataset are in Appendix B.2.
@@ 5.2 Performance Comparisons
Compared Methods. We compare our method with three classes of baselines: (1) Traditional sequential recommendation methods, i.e., GRU4Rec (Hidasi et al., 2015), Caser (Tang & Wang, 2018), HGN (Ma et al., 2019), BERT4Rec (Sun et al., 2019), SASRec (Kang & McAuley, 2018) and LightSANs (Fan et al., 2021). (2) Self-supervised sequential recommendation methods, i.e., S3-Rec (Zhou et al., 2020), DuoRec (Qiu et al., 2022) and MAERec (Ye et al., 2023). (3) G-LLMRec method: Open-P5LLaMa [^5] For Open-P5, we adopt the version of LLaMa as the foundation model in their code repository implementation to ensure the best results are achieved. (Xu et al., 2023a). (4) E-LLMRec method: E4SRec (Li et al., 2023a). A detailed introduction to these baselines can be found in Appendix B.3. It should be noted that we did not select various G-LLMRec methods or E-LLMRec methods as baselines. This is because the differences between each LLM-based method are minimal, and our model is a universal approach that is not confined to a specific model type. Our primary focus is to improve the efficiency of language model utilization. Hence, we opted to select one G-LLMRec method (Open-P5) and one E-LLMRec method (E4SRec) as baselines.
Table 2: Experimental results (%) on the Cloth and Movie dataset. The missing MRR value of Open-P5 is unavailable due to the time complexity constrictions. The number on the left of the arrow is the layers $N$ of the student model. The left number on the right of the arrow is the layers $M$ of the teacher model. For Open-P5, we adopt LLaMa as their backbone. We highlight the methods with the best and second-best average performances. We also give the average ranking of each evaluation metric. Moreover, E4SRec4, which has the same number of layers as our SLMRec, is also marked.
Model Cloth Movie Rank
HR@1 HR@5 NDCG@5 MRR HR@1 HR@5 NDCG@5 MRR
Caser 9.66 15.18 12.66 13.03 4.27 14.96 9.57 10.36 13.50
GRU4Rec 13.79 15.46 14.64 15.15 10.56 19.47 15.11 15.46 9.25
BERT4Rec 13.60 14.66 14.14 14.59 9.68 14.91 12.40 12.74 11.63
SASRec 13.08 16.94 15.01 15.76 5.57 16.80 11.17 12.08 11.63
HGN 15.96 18.70 17.30 18.27 7.54 19.20 13.42 14.73 6.50
LightSANs 14.12 20.32 17.30 16.86 6.08 17.54 11.81 12.82 8.00
S3-Rec 14.10 18.67 16.10 16.95 7.75 20.39 15.69 14.34 7.50
DuoRec 13.06 18.29 15.79 15.42 10.07 20.37 17.96 16.61 7.88
MAERec 13.29 18.35 15.68 16.13 8.89 20.24 16.03 15.28 8.38
Open-P5 14.13 17.68 17.02 - 12.66 21.98 17.13 - 5.67
E4SRec 16.71 19.45 18.09 18.77 14.74 23.79 19.45 19.74 1.75
E4SRec8 15.30 18.54 16.91 17.60 13.32 22.49 17.99 18.46 4.00
E4SRec4 14.58 18.05 16.32 17.01 11.80 21.54 16.73 17.20 5.75
$\mathrm{SLMRec}_{4\leftarrow 8}$ 16.69 19.47 18.07 18.74 15.29 24.25 19.90 20.36 1.50
@@
Model Music Sport Rank
HR@1 HR@5 NDCG@5 MRR HR@1 HR@5 NDCG@5 MRR
Caser 0.71 3.28 1.96 2.29 1.05 3.75 2.39 2.84 13.50
GRU4Rec 1.89 3.22 2.57 3.08 5.26 7.75 6.52 7.08 10.13
BERT4Rec 2.10 3.16 2.64 3.11 4.81 6.70 5.79 6.26 10.63
SASRec 1.82 5.72 3.79 4.51 4.70 8.43 6.59 7.24 8.75
HGN 2.01 5.49 3.82 4.17 3.42 6.24 4.83 5.30 10.50
LightSANs 1.05 4.06 2.54 3.00 5.18 8.94 7.07 7.72 8.25
S3-Rec 2.48 7.37 4.94 4.68 4.14 8.49 6.89 7.35 6.88
DuoRec 1.84 4.50 3.19 3.04 4.13 8.81 7.03 6.64 9.13
MAERec 2.19 6.35 4.67 3.96 4.01 8.35 6.65 6.98 8.63
Open-P5 4.35 8.12 6.74 - 5.49 8.50 6.92 - 5.33
E4SRec 5.62 9.29 7.50 7.98 6.40 9.67 8.05 8.70 1.75
E4SRec8 5.46 8.86 7.21 7.74 5.48 8.63 7.06 7.76 3.63
E4SRec4 5.33 8.75 7.08 7.59 5.41 8.65 7.04 7.72 4.50
$\mathrm{SLMRec}_{4\leftarrow 8}$ 5.72 9.15 7.48 8.03 6.62 9.83 8.25 8.89 1.25
@@
Quantitative Results (RQ1). Tables 23 showcase the quantitative comparison of four large-scale sequential recommendation datasets. From our analysis, we have several insightful observations: (1) LLM-based recommendation methods exhibit substantial improvements over traditional sequential recommendation (TSR) methods, primarily due to their enhanced modeling capacity which adeptly extracts informative sequential interest patterns. (2) Our model, $\mathrm{SLMRec}_{4\leftarrow 8}$, outperforms the teacher model E4SRec8 by leveraging knowledge distillation within the hidden layers. By refraining from applying this constraint prior to the prediction phase, we enable the final representation to organically gravitate towards the label—yielding an approximate 8% enhancement in performance in comparison to the teacher model. (3) Introducing vanilla knowledge distillation techniques into LLMRec, without altering the model structure, allows $\mathrm{SLMRec}_{4\leftarrow 8}$ to achieve a marginally superior performance compared to E4SRec32. This suggests that small language models equipped with efficacious training strategies can rival, or even exceed, larger language models in the sequential recommendation task. This phenomonon is also matched with our therotical justification in Section 6.
@@
Model Efficiency (RQ2). We report the time efficiency and parameters of comparative baselines and our model in Table 5. All time and parameter metrics represent the average across the four datasets reported. Inference time evaluates the prediction ranking among 1,000 candidate items for each user. Detailed training and inference times for each dataset are provided in Appendix B.4. The Open-P5, an LLMRec model based on generative methods, offers a reasonable training duration. Yet, during the inference phase, it becomes considerably time-consuming (4942 hours) as it necessitates generating a substantial pool of candidate items (for instance, 1000). Owing to the intrinsic workings of generative LLMs, employing generation-based LLMRec models for the comprehensive ranking of extensive item sets is not advised. Our model outperforms E4SRec with enhanced efficiency, maintaining only 13% and 14% in E4SRec’s parameters for training and inference, respectively. Moreover, our SLMRec demonstrates a remarkable gain in speed, being 6.6 times faster during training and 8.0 times quicker in inference than E4SRec.
@@
Ablation Study (RQ3). As shown in Table 4, SLMRec, when enhanced with various knowledge regularizers (namely $\mathcal{D}_{cos},\mathcal{D}_{norm}$ and $\mathcal{L}_{ms}$), demonstrates improved performance. The regularizers $\mathcal{D}_{cos}$ and $\mathcal{D}_{norm}$ aid SLMRec in aligning its intermediate representations with those of the teacher model, thereby endowing it with more potent representational extraction capabilities. Meanwhile, $\mathcal{L}_{ms}$ steers the model to assimilate domain knowledge pertinent to recommendation systems within its preliminary layers. The ablation study in Music and Sport domain can be found in Appendix B.5.
Table 4: Experiment results (%) of ablation study.
SLMRec Cloth Movie
HR@1 HR@5 NDCG@5 MRR HR@1 HR@5 NDCG@5 MRR
+$\mathcal{D}_{cos}$ 16.10 18.85 17.48 18.17 14.83 23.08 19.08 19.45
+$\mathcal{D}_{cos}$,$\mathcal{D}_{norm}$ 16.28 19.12 17.69 18.40 14.86 23.89 19.36 19.84
+$\mathcal{D}_{cos}$,$\mathcal{L}_{ms}$ 16.85 19.05 17.96 18.59 15.05 23.48 19.40 19.76
+$\mathcal{D}_{cos}$,$\mathcal{D}_{norm}$,$\mathcal{L}_{ms}$ 16.69 19.47 18.07 18.74 15.29 24.25 19.90 20.36
@@ 5.3 Model Study
Figure 4: Experiment results (%) of online KD and block number $B$ in the Cloth dataset.
6 Theoretical Justifications
Beyond empirical experiments, we aim to provide insights into why small language models can perform as effectively as large language models in learning desirable user representations. Specifically, we focus on the feature propagation process within a single layer of an LLM, as outlined below:
$\mathbf{H}^{(k)}=\mathbf{H}^{(k-1)}+\mathbf{A}^{(k)}\mathbf{H}^{(k-1)},$ (7)
where $\mathbf{H}^{(k)}$ represents the hidden representation of the $k$-th layer, and $\mathbf{A}^{(k)}$ is the attention matrix. In LLaMa, the attention matrix is defined as $\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q^{\prime}}\mathbf{K^{\prime}}^{\top}}{\sqrt{d_{k}}}\right)$, where $\mathbf{Q^{\prime}}$ and $\mathbf{K^{\prime}}$ incorporate rotational encoding (Su et al., 2024). Our analysis is hinged on interpreting the stack of propagation layers of Transformers as optimization dynamics for minimizing energies of certain forms (Shuman et al., 2013; Kalofolias, 2016; Fu et al., 2022; Wu et al., 2024).
Proposition 1.
Given the updating matrix $\hat{\mathbf{A}}^{(k)}=\mathbf{A}^{(k)}+\mathbf{I}$, Eqn. 7 is equivalent to a gradient descent step with respect to the following optimization problem:
$\min_{H}\left\|\mathbf{H}-\hat{\mathbf{A}}^{(k)}\mathbf{H}^{(k-1)}\right\|_{2}^{2}$ (8)
As $\mathbf{A}^{(k)}$ changes across layers, multi-layer attention models can be interpreted as a series of iterative descent steps, each focusing on layer-specific denoising objectives. We will show that this multi-layer structure can be simplified into a single-layer model while retaining the same denoising effectiveness.
Proposition 2.
For any $K$-layer attention model (where $K$ is an arbitrary positive integer) with the layer-wise updating rule defined by Eqn. 7, there exists $\mathbf{C}^{*}$ such that one gradient descent step for the optimization problem (from the initial embeddings $\mathbf{H}^{(0)}$)
$\min_{\mathbf{H}}\left\|\mathbf{H}-\mathbf{C}^{*}\mathbf{H}^{(0)}\right\|_{2}^{2},$ (9)
where $\mathbf{C}^{*}$ associated with $\mathbf{A}$, can yield the output embeddings $\mathbf{H}^{(K)}$ of the $K$-layer model.
These findings indicate that for any multi-layer stacked decoder, an equivalent single-layer decoder can be constructed to encode hidden representations in a similar way. Moreover, while the multi-layer model optimizes distinct objectives at each layer, this may introduce redundancy when compared to a single-layer model that achieves its objective in a single step. Consistent with the motivation underlying our framework design, we employ knowledge distillation (KD) to guide the one-layer network, enabling it to streamline the learning process and replicate the feature extraction capabilities of a multi-layer network.
@@ 7 Related Work
In this section, we introduce the most related background and scientific investigations to this work, which are roughly divided into five categories, i.e., 1) Sequential Recommendation, 2) Knowledge Distillation (KD), 3) Depth-wise Knowledge of LLMs, 4) Model Pruning, and 5) Parameter-Efficient Fine-Tuning (PEFT). For details on sections three through five, please refer to Appendix C.
@@
Sequential Recommendation. Traditional Sequential Recommendation (TSR) methods (Wu et al., 2017; Hidasi et al., 2015; Kang & McAuley, 2018) primarily focus on developing various temporal encoders to capture short- and long-term user interests. The evolution of temporal sequential encoders has progressed from LSTM units (Wu et al., 2017) and GRU units (Hidasi et al., 2015), to more advanced architectures such as graph neural networks (He et al., 2020; Xu et al., 2023b; 2024c), self-attention layers (Kang & McAuley, 2018; Xu et al., 2024b), and Transformer models (Sun et al., 2019). Following the triumph of large language models (LLMs), researchers have begun leveraging open-source LLMs (Touvron et al., 2023) to construct their recommendation systems (Ji et al., 2024; Bao et al., 2023; Wei et al., 2024). G-LLMRec methods (Geng et al., 2022; Xu et al., 2023a; Zhang et al., 2023b; Liao et al., 2023; Mei & Zhang, 2023) generate the next item based on historical sequences, while E-LLMRec approaches (Li et al., 2023a; Zhu et al., 2023; Wang et al., 2024) use LLMs as feature extractors to learn user representations for prediction. More recently,  (Zhai et al., 2024) introduces a generative sequential framework scalable up to GPT-3 dimensions. LLM-based recommendation systems frequently outperform TSR models by a margin of 20% (Li et al., 2023a; Liao et al., 2023; Wang et al., 2024), also increasing the parameters by nearly 100 times compared to TSR models. Therefore, the deployment of LLMRec models in real-world platforms is heavily constrained by computational resources.
@@
Knowledge Distillation (KD). Training a smaller “student" model on the distribution predicted by a large “teacher" model is known as a powerful knowledge distillation technique (Hinton et al., 2015). The fundamental insight behind this is to transform the knowledge and capabilities of the teacher into more compact, compressed, and possibly skill-specific representations (Jiao et al., 2020; Gu et al., 2024). For those cases when the student only has access to the output tokens generated by the teacher, another way of KD is data distillation (Eldan & Li, 2023; Li et al., 2023b; Fu et al., 2023; Hsieh et al., 2023). This technique first generates high-quality synthetic data by prompting the larger teacher model. The synthetic data are then used to enhance the student’s capabilities by fine-tuning.
8 Conclusions